Processes and Threads

What are processes?

How does the operating system manage them?

Nick Urbanik nicku@nicku.org

A computing department

Copyright Conditions: Open Publication License

(see http://www.opencontent.org/openpub/)

Introduction What is a process?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 2 What is a process? — 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 3 What is a thread? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 4 Program counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 5 Environment of a process . . . . . . . . . . . . . . . . . . . . . . . . . . slide 6 Permissions of a Process. . . . . . . . . . . . . . . . . . . . . . . . . . . slide 7

Multitasking Multitasking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 8 Multitasking — 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 9 Multitasking — 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 10

Start of Process Birth of a Process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 11 Process tree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 12

Scheduler Scheduler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 13 When to Switch Processes?. . . . . . . . . . . . . . . . . . . . . . . . slide 14 Scheduling statistics: vmstat . . . . . . . . . . . . . . . . . . . . . . slide 15 Interrupts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 16

Process States Process States . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 17 What is Most Common State? . . . . . . . . . . . . . . . . . . . . . slide 18 Most Processes are Blocked . . . . . . . . . . . . . . . . . . . . . . . slide 19 Linux Process States . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 20 Linux Process States — 2. . . . . . . . . . . . . . . . . . . . . . . . . slide 21 Linux Process States — 3. . . . . . . . . . . . . . . . . . . . . . . . . slide 22 Process States: vmstat . . . . . . . . . . . . . . . . . . . . . . . . . . slide 23 Tools for monitoring processes . . . . . . . . . . . . . . . . . . . . . slide 24 Monitoring processes in Win 2000 . . . . . . . . . . . . . . . . . . . slide 25

top Process Monitoring — top . . . . . . . . . . . . . . . . . . . . . . . . slide 27 load average . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 28 top: process states. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 29 top and memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 30 Virtual Memory: suspended processes . . . . . . . . . . . . . . . . slide 31 Suspended Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 32

Process Control Blocks OS Process Control Structures . . . . . . . . . . . . . . . . . . . . . slide 34 What is in a PCB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 35 Context Switch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 36 Execution Context. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 37 Program Counter in PCB. . . . . . . . . . . . . . . . . . . . . . . . . slide 38 PCB Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 39 PCB Example Diagram . . . . . . . . . . . . . . . . . . . . . . . . . . slide 40 PCB Example — Continued . . . . . . . . . . . . . . . . . . . . . . . slide 41 Address of I/O instructions . . . . . . . . . . . . . . . . . . . . . . . slide 42

System Calls System Calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 44 File I/O system calls: a sidetrack . . . . . . . . . . . . . . . . . . . slide 45 init. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 46 SUID, SGID and IDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 47 Other system calls: getting process info . . . . . . . . . . . . . . . slide 48 fork(): what it does . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 49 Using fork(): pseudocode . . . . . . . . . . . . . . . . . . . . . . . . slide 50 Simple fork() Example (no Checking) . . . . . . . . . . . . . . . slide 51 An example using fork(). . . . . . . . . . . . . . . . . . . . . . . . . slide 52 Example using fork()—(contd.). . . . . . . . . . . . . . . . . . . . slide 53 Output of fork-example.c: . . . . . . . . . . . . . . . . . . . . . . . slide 54 Running fork-example again . . . . . . . . . . . . . . . . . . . . . . slide 55 Why two “before fork” messages? . . . . . . . . . . . . . . . . . slide 56 So what does this show?. . . . . . . . . . . . . . . . . . . . . . . . . . slide 57 Running another program — exec(). . . . . . . . . . . . . . . . . slide 58 execve() system call . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 59 fork() — exec() Example . . . . . . . . . . . . . . . . . . . . . . . slide 60 Using execl() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 61 print.c: a program we call . . . . . . . . . . . . . . . . . . . . . . . slide 62 Calling ./print using execl(). . . . . . . . . . . . . . . . . . . . . slide 63 vfork() sytem call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 64 wait(), waitpid() system calls . . . . . . . . . . . . . . . . . . . . slide 65 wait(), waitpid() system calls . . . . . . . . . . . . . . . . . . . . slide 66

A shell program Part of Simple Shell Program . . . . . . . . . . . . . . . . . . . . . . slide 67 Windows and Processes . . . . . . . . . . . . . . . . . . . . . . . . . . slide 68 Windows and Processes — 2. . . . . . . . . . . . . . . . . . . . . . . slide 69 CreateProcess() prototype. . . . . . . . . . . . . . . . . . . . . . . slide 70 CreateProcess() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 71 Example: CreateProcess() . . . . . . . . . . . . . . . . . . . . . . . slide 72 Processes in Linux, Unix, Windows . . . . . . . . . . . . . . . . . . slide 73

IPC Problem with Processes . . . . . . . . . . . . . . . . . . . . . . . . . . slide 75 Interprocess Communication (IPC) . . . . . . . . . . . . . . . . . . slide 76 IPC — Shared Memory . . . . . . . . . . . . . . . . . . . . . . . . . . slide 77 IPC — Signals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 78 Signals and the Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 79

Threads Threads and Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 81 Threads have own. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 82 Threads share a lot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 83 Threads in Linux, Unix . . . . . . . . . . . . . . . . . . . . . . . . . . slide 84 hello.c: a simple threaded program. . . . . . . . . . . . . . . . . slide 85 Compile POSIX Threads . . . . . . . . . . . . . . . . . . . . . . . . . slide 86 pthread create(). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 87 pthread create(). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 88 Problem with threads: . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 89

Race Condition Race Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 91 Critical Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 92 Race Condition — one possibility . . . . . . . . . . . . . . . . . . . slide 93 Example — another possibility . . . . . . . . . . . . . . . . . . . . . slide 94 Solution: Synchronisation. . . . . . . . . . . . . . . . . . . . . . . . . slide 95 File Locking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 96

Synchronisation Synchronisation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 98 Semaphores. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 99 Semaphores — 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 100 POSIX and Win32. . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 101 Threads Example 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 102 Threads Example 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 103 Threads Example 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 104 Condition Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 105 Monitors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 106

Spinlocks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 107 Summary and References

Summary — Process States, Scheduling. . . . . . . . . . . . . . slide 109 Summary — Processes and Threads . . . . . . . . . . . . . . . . slide 110 Summary — Synchronisation . . . . . . . . . . . . . . . . . . . . . slide 111 References. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . slide 112

What is a process?

• A process is a program in execution

• Each process has a process ID

• In Linux,

$ ps ax

• prints one line for each process.

• A program can be executed a number of times simultaneously.

∘ Each is a separate process.

OSSI — ver. 1.5 Processes — slide 2

What is a process? — 2

• A process includes current values of:

∘ Program counter ∘ Registers ∘ Variables

• A process also has:

∘ The program code ∘ It’s own address space, independent of other processes ∘ A user that owns it ∘ A group owner ∘ An environment and a command line

• This information is stored in a process control block, or task de-

scriptor or process descriptor ∘ a data structure in the os, in the process table ∘ See slides starting at §34.

OSSI — ver. 1.5 Processes — slide 3

What is a thread?

• A thread is a lightweight process ∘ Takes less cpu power to start, stop

• Part of a single process

• Shares address space with other threads in the same process

• Threads can share data more easily than processes

• Sharing data requires synchronisation, i.e., locking — see slide 95.

• This shared memory space can lead to complications in program-

ming:

“Threads often prevent abstraction. In order to prevent deadlock. you often need to know how and if the library you are using uses threads in order to avoid deadlock problems. Similarly, the use of threads in a library could be affected by the use of threads at the application layer.” – David Korn

See page 180, ESR in references, §112.

OSSI — ver. 1.5 Processes — slide 4

Program counter

• The code of a process occupies memory

• The Program counter (PC) is a cpu register

• PC holds a memory address…

• … of the next instruction to be fetched and executed

OSSI — ver. 1.5 Processes — slide 5

Environment of a process

• The environment is a set of names and values

• Examples:

  PATH=/usr/bin:/bin:/usr/X11R6/bin
  HOME=/home/nicku
  SHELL=/bin/bash

• In Linux shell, can see environment by typing:

  $ set

OSSI — ver. 1.5 Processes — slide 6

Permissions of a Process

• A process executes with the permissions of its owner ∘ The owner is the user that starts the process

• A Linux process can execute with permissions of another user or group

• If it executes as the owner of the program instead of the owner of the process, it is called set user id

• Similarly for set group id programs

OSSI — ver. 1.5 Processes — slide 7

Multitasking

• Our lab PCs have one main cpu

∘ But multiprocessor machines are becoming increasingly com-

mon ∘ Linux 2.6.x kernel scales to 16 cpus

• How execute many processes “at the same time”?

OSSI — ver. 1.5 Processes — slide 8

Multitasking — 2

• CPU rapidly switches between processes that are “ready to run”

• Really: only one process runs at a time

• Change of process called a context switch ∘ See slide §36

• With Linux: see how many context switches/second using vmstat

under “system” in column “cs”

OSSI — ver. 1.5 Processes — slide 9

Multitasking — 3

• This diagram shows how the scheduler gives a “turn” on the cpu to each of four processes that are ready to run

Birth of a Process

• In Linux, a process is born from a fork() system call

∘ A system call is a function call to an operating system service

provided by the kernel

• Each process has a parent

• The parent process calls fork()

• The child inherits (but cannot change) the parent environment, open files

• Child is identical to parent, except for return value of fork(). ∘ Parent gets child’s process ID (pid) ∘ Child gets 0

OSSI — ver. 1.5 Processes — slide 11

Process tree

• Processes may have parents and children

• Gives a family tree

• In Linux, see this with commands:

  $ pstree

or

  $ ps axf

OSSI — ver. 1.5 Processes — slide 12

Scheduler

• OS decides when to run each process that is ready to run (“runable”)

• The part of OS that decides this is the scheduler

• Scheduler aims to:

∘ Maximise CPU usage ∘ Maximise process completion ∘ Minimise process execution time ∘ Minimise waiting time for ready processes ∘ Minimise response time

OSSI — ver. 1.5 Processes — slide 13

When to Switch Processes?

• The scheduler may change a process between executing (or run-

ning) and ready to run when any of these events happen: ∘ clock interrupt ∘ I/O interrupt ∘ Memory fault ∘ trap caused by error or exception ∘ system call

• See slide §17 showing the running and ready to run process states.

OSSI — ver. 1.5 Processes — slide 14

Scheduling statistics: vmstat

• The “system” columns give statistics about scheduling: ∘ “cs” — number of context switches per second ∘ “in” — number of interrupts per second

• See slide §36, man vmstat

OSSI — ver. 1.5 Processes — slide 15

Interrupts

• Will discuss interrupts in more detail when we cover I/O

• An interrupt is an event (usually) caused by hardware that causes: ∘ Saving some CPU registers ∘ Execution of interrupt handler ∘ Restoration of CPU registers

• An opportunity for scheduling

OSSI — ver. 1.5 Processes — slide 16

Process States

OSSI — ver. 1.5 Processes — slide 17

What is Most Common State?

• Now, my computer has 160 processes.

• How many are running, how many are ready to run, how many are blocked?

• What do you expect is most common state?

OSSI — ver. 1.5 Processes — slide 18

Most Processes are Blocked

9:41am  up 44 days, 20:12,  1 user,  load average: 2.02, 2.06, 2.13
160 processes: 145 sleeping, 2 running, 13 zombie, 0 stopped

• Here you see that most are sleeping, waiting for input!

• Most processes are “I/O bound”; they spend most time waiting for input or waiting for output to complete

• With one cpu, only one process can actually be running at one time

• However, surprisingly few processes are ready to run

• The load average is the average number of processes that are in the ready to run state.

• In output from the top program above, see over last 60 seconds,

there are 2.02 processes on average in rtr state

OSSI — ver. 1.5 Processes — slide 19

Linux Process States Linux Process States — 2

• Running — actually contains two states:

∘ executing, or ∘ ready to execute

• Interruptable — a blocked state

∘ waiting for event, such as:

– end of an I/O operation,

– availability of a resource, or

– a signal from another process

• Uninterruptable — another blocked state ∘ waiting directly on hardware conditions ∘ will not accept any signals (even SIGKILL)

OSSI — ver. 1.5 Processes — slide 21

Linux Process States — 3

• Stopped — process is halted ∘ can be restarted by another process ∘ e.g., a debugger can put a process into stopped state

• Zombie — a process has terminated

∘ but parent did not wait() for it (see slide 33)

OSSI — ver. 1.5 Processes — slide 22

Process States: vmstat

• The “procs” columns give info about process states:

• “r” — number of processes that are in the ready to run state • “b” — number of processes that are in the uninterruptable blocked

state

OSSI — ver. 1.5 Processes — slide 23

Tools for monitoring processes

• Linux provides:

• vmstat

∘ Good to monitor over time:

      $ vmstat 5
• procinfo

∘ Easier to understand than vmstat ∘ Monitor over time with

      $ procinfo -f

• View processes with top — see slides 27 to §30

• The system monitor sar shows data collected over time: See

  man sar; investigate sar -c and sar -q

• See the utilities in the procps software package. You can list them with

  $ rpm -ql procps

  ps free pgrep      pkill pmap skill      slabtop snice tload top uptime vmstat     {w \atop watch}

OSSI — ver. 1.5 Processes — slide 24

Monitoring processes in Win 2000

• Windows 2000 provides a tool:

• Start → Administrative Tools → Performance.

• Can use this to monitor various statistics

OSSI — ver. 1.5 Processes — slide 25

Process Monitoring with top

Process Monitoring — top

   08:12:13  up 1 day, 13:34,  8 users,  load average: 0.16, 0.24, 0.49
111 processes: 109 sleeping, 1 running, 1 zombie, 0 stopped
CPU states:  cpu    user    nice  system    irq  softirq  iowait    idle
           total    0.0%    0.0%    3.8%   0.0%     0.0%    0.0%   96.1%
Mem:   255608k av,  245064k used,   10544k free,       0k shrd,   17044k buff
       152460k active,              63236k inactive
Swap: 1024120k av,  144800k used,  879320k free                  122560k cached
  PID USER     PRI  NI  SIZE  RSS SHARE STAT %CPU %MEM   TIME CPU COMMAND
 1253 root      15   0 73996  13M 11108 S     2.9  5.5  19:09   0 X
 1769 nicku     16   0  2352 1588  1488 S     1.9  0.6   2:10   0 magicdev
23548 nicku     16   0  1256 1256   916 R     1.9  0.4   0:00   0 top
    1 root      16   0   496  468   440 S     0.0  0.1   0:05   0 init
    2 root      15   0     0    0     0 SW    0.0  0.0   0:00   0 keventd
    3 root      15   0     0    0     0 SW    0.0  0.0   0:00   0 kapmd
    4 root      34  19     0    0     0 SWN   0.0  0.0   0:00   0 ksoftirqd/0
    6 root      15   0     0    0     0 SW    0.0  0.0   0:00   0 bdflush
    5 root      15   0     0    0     0 SW    0.0  0.0   0:11   0 kswapd

OSSI — ver. 1.5 Processes — slide 27

top: load average

08:12:13  up 1 day, 13:34,  8 users,  load average: 0.16, 0.24, 0.49

• load average is measured over the last minute, five minutes, fifteen minutes

• Over that time is the average number of processes that are ready to run, but which are not executing

• A measure of how “busy” a computer is.

OSSI — ver. 1.5 Processes — slide 28

top: process states

111 processes: 109 sleeping, 1 running, 1 zombie, 0 stopped

sleeping Most processes (109/111) are sleeping, waiting for i^{/}o

running This is the number of processes that are both ready to run and are executing

zombie There is one process here that has terminated, but its parent did not wait() for it.

• The wait() system calls are made by a parent process, to get the exit() status of its child(ren).

• This call removes the process control block from the process table, and the child process does not exist any more. (§34)

stopped When you press [figure: Control-z]in a shell, you will increase this

number by 1

OSSI — ver. 1.5 Processes — slide 29

top: Processes and Memory

PID USER     PRI  NI  SIZE  RSS SHARE STAT %CPU %MEM   TIME CPU COMMAND
1253 root      15   0 73996  13M 11108 S     2.9  5.5  19:09   0 X

SIZE This column is the total size of the process, including the part which is swapped (paged out) out to the swap partition or swap file

Here we see that the process X uses a total of 73,996 Kb, i.e., 73,996 × 1024 bytes ≈ 72MB, where here 1MB = 2^{20} bytes.

RSS The resident set size is the total amount of ram that a process uses, including memory shared with other processes. Here X uses a total of 13MB ram, including ram shared with other processes.

SHARE The amount of shared memory is the amount of ram that this process shares with other processes. Here X shares 11,108 KB with other processes.

We can see that the total amount of ram used exclusively by one process is rss − share. Here we see that X uses about 13 × 2^{20} − 11,108 × 2^{10} ≈ 2 MB

OSSI — ver. 1.5 Processes — slide 30

Virtual Memory: suspended processes

• With memory fully occupied by processes, could have all in blocked state!

• cpu could be completely idle, but other processes waiting for ram

• Solution: virtual memory

∘ will discuss details of vm in memory management lecture

• Part or all of process may be saved to swap partition or swap file

OSSI — ver. 1.5 Processes — slide 31

Suspended Processes

• Could add more states to process state table:

∘ ready and suspended ∘ blocked and suspended

OSSI — ver. 1.5 Processes — slide 32

Process Control Blocks

The Process Table

Data structure in OS to hold information about a process

OS Process Control Structures

• Every os provides process tables to manage processes

• In this table, the entries are called process control blocks (pcbs), process descriptors or task descriptors. We will use the abbreviation pcb.

• There is one pcb for each process

• in Linux, pcb is called task struct, defined in

  include/linux/sched.h

∘ In a Fedora Core or Red Hat system, you will find it in the file

      /usr/src/linux-2.*/include/linux/sched.h if you have

installed the kernel-source software package

OSSI — ver. 1.5 Processes — slide 34

What is in a PCB

• In slide §3, we saw that a pcb contains:

∘ a process id (pid) ∘ process state (i.e., executing, ready to run, sleeping waiting

for input, stopped, zombie) ∘ program counter, the cpu register that holds the address of

the next instruction to be fetched and executed ∘ The value of other CPU registers the last time the program

was switched out of executing by a context switch — see slide §36

∘ scheduling priority ∘ the user that owns the process ∘ the group that owns the process ∘ pointers to the parent process, and child processes ∘ Location of process’s data and program code in memory ∘ List of allocated resources (including open files)

• pcb holds the values as they were when process was last switched

out of executing by a context switch — see slide §36

OSSI — ver. 1.5 Processes — slide 35

Context Switch

• os does a context switch when:

∘ stop current process from executing, and

∘ start the next ready to run process executing on cpu

• os saves the execution context (see §37) to its pcb

• os loads the ready process’s execution context from its pcb

• When does a context switch occur?

∘ When a process blocks, i.e., goes to sleep, waiting for input

or output (i^{/}o), or ∘ When the scheduler decides the process has had its turn of the cpu, and it’s time to schedule another ready-to-run process

• A context switch must be as fast as possible, or multitasking will

be too slow

∘ Very fast in Linux os

OSSI — ver. 1.5 Processes — slide 36

Execution Context

• Also called state of the process (but since this term has two meanings, we avoid that term here), process context or just context

• The execution context is all the data that the os must save to stop one process from executing on a cpu, and load to start the next process running on a cpu

• This includes the content of all the cpu registers, the location of

the code, …

∘ Includes most of the contents of the process’s pcb.

OSSI — ver. 1.5 Processes — slide 37

Program Counter in PCB

• What value is in the program counter in the pcb?

• If it is not executing on the cpu,

∘ The address of the next cpu instruction that will be fetched and executed the next time the program starts executing

• If it is executing on the cpu,

∘ The address of the first cpu instruction that was fetched and executed when the process began executing at the last context switch (§36)

OSSI — ver. 1.5 Processes — slide 38

Process Control Blocks—Example

• The diagram in slide §40 shows three processes and their process control blocks.

• There are seven snapshots t_{0}, t_{1}, t_{2}, t_{3}, t_{4}, t_{5} and t_{6} at which the scheduler has changed process (there has been a context switch— §36)

• On this particular example cpu, all i^{/}o instructions are 2 bytes long

• The diagram also shows the queue of processes in the:

∘ Ready queue (processes that are ready to run, but do not have a cpu to execute on yet)

∘ Blocked, or Wait queue, where the processes have been

blocked because they are waiting for i^{/}o to finish.

OSSI — ver. 1.5 Processes — slide 39

PCB Example: Diagram

PSfrag replacements

t_{0} t_{1} t_{2} t_{3} t_{4} t_{5} t_{6}

OSSI — ver. 1.5 Processes — slide 40

PCB Example — Continued

• In slide §40,

∘ The times t_{0}, t_{1}, t_{2}, t_{3}, t_{4}, t_{5} and t_{6} are when the scheduler

has selected another process to run. ∘ Note that these time intervals are not equal, they are just the

points at which a scheduling change has occurred.

• Each process has stopped at one stage to perform i^{/}o

∘ That is why each one is put on the wait queue once during

its execution.

• Each process has performed i^{/}o once

OSSI — ver. 1.5 Processes — slide 41

What is the address of I/O instructions?

• We are given that all i^{/}o instructions in this particular example are two bytes long (slide §39)

∘ We can see that when the process is sleeping (i.e., blocked), then the program counter points to the instruction after the i^{/}o instruction

∘ So for process P1, which blocks with program counter pc = C0DE_{16}, the i^{/}o instruction is at address C0DE_{16} − 2 = C0DC_{16}

∘ for process P2, which blocks with program counter pc = FEED_{16}, the i^{/}o instruction is at address FEED_{16} −2 = FEEB_{16}

∘ for process P3, which blocks with program counter pc = D1CE_{16}, the i^{/}o instruction is at address D1CE_{16} − 2 = D1CC_{16}

OSSI — ver. 1.5 Processes — slide 42

Process System Calls

How the OS controls processes How you use the OS to control processe

Major process Control System Calls

• fork() — start a new process

• execve() — replace calling process with machine code from another program file

• wait(), waitpid() — parent process gets status of its’ child after the child has terminated, and cleans up the process table entry for the child (stops it being a zombie)

• exit() — terminate the current process

OSSI — ver. 1.5 Processes — slide 44

File I/O system calls: a sidetrack

  #include <unistd.h>
  ssize_t read( int filedes, void *buf,
                size_t nbytes );

• returns number of bytes read, 0 at end of file, −1 on error

  ssize_t write( int filedes, void *buf,
                 size_t nbytes );

• returns number of bytes written, else −1 on error

• Note: these are unbuffered, that is, they have effect “immediately”.

• This is different from stdio.h functions, which are buffered for

efficiency.

OSSI — ver. 1.5 Processes — slide 45

Process IDs and init

• Every process has a process id (pid)

• process 0 is the scheduler, part of kernel

• process 1 is init, the parent of all other processes ∘ a normal user process, not part of kernel ∘ program file is /sbin/init

• All other processes result from init calling the fork() system call

• This is the only way a new process is created by the kernel

OSSI — ver. 1.5 Processes — slide 46

SUID, SGID and IDs

• Every process has six or more ids associated with it

• uid and gid of person who executes program file:

∘ real user id, real group id

• ids used to calculate permissions:

∘ Effective uid, Effective gid

• ids saved when use exec() system call:

∘ Saved set-user-ID, saved set-group-ID

∘ idea is can drop special privileges and return to executing with real uid and real gid when privilege is no longer required

OSSI — ver. 1.5 Processes — slide 47

Other system calls: getting process info

  #include <sys/types.h>
  #include <unistd.h>

• pid t getpid(void); returns PID of calling process

• pid t getppid(void); returns PID of parent

• uid t getuid(void); returns real user ID of process

• uid t geteuid(void); returns effective user ID of process

• gid t getgid(void); returns real group ID of process

• gid t getegid(void); returns effective group ID of process

OSSI — ver. 1.5 Processes — slide 48

fork(): what it does

  #include <sys/types.h>
  #include <unistd.h>
  pid_t fork(void);

• returns 0 in child

• returns pid of child in parent

• returns −1 if error

OSSI — ver. 1.5 Processes — slide 49

Using fork(): pseudocode

if ( ( pid = fork() ) < 0 )

fork error has happened

else if ( pid == 0 ) /∗ I am the child ∗/

do things the child process should do

else /∗ I am the parent ∗/ do things the parent should do

OSSI — ver. 1.5 Processes — slide 50

Simple fork() Example (no Checking)

#include <stdio.h>
#include <unistd.h>

int main() {

int pid = fork();

printf( "PID is %d\n", pid ); if ( pid == 0 )

               printf( "I'm the child\n" );

else

               printf( "I'm the parent\n" );

}

An example using fork()

#include <sys/types.h>
#include <unistd.h>
#include <stdio.h>

int glob = 6;

char buf[] = "a write to standard output\n";

int main( void ) {

int var = 88; /∗ local variable on the stack ∗/ pid t pid; if ( write( STDOUT FILENO, buf, sizeof (buf) − 1 )

≠ sizeof( buf ) − 1 ) { fprintf( stderr, "write error" ); exit( 1 );

}

OSSI — ver. 1.5 Processes — slide 52

Example using fork()—(contd.)

printf( "before fork\n" );

if ( ( pid = fork() ) < 0 ) { fprintf( stderr, "fork error\n" ); exit( 1 );

} else if ( pid == 0 ) { /∗ child ∗/ ++glob; ++var;

} else

sleep( 2 ); /∗ parent ∗/

printf( "pid = %d, glob = %d, var = %d\n",

getpid(), glob, var ); exit( 0 );

}

Output of fork-example.c:
$ gcc -o fork-example fork-example.c
$ ./fork-example
a write to standard output
before fork
pid = 7118, global = 7, var = 89 child's vars changed
pid = 7117, global = 6, var = 88 parent's copy not changed

OSSI — ver. 1.5 Processes — slide 54

Running fork-example again
$ ./fork-example > tmp.out
$ cat tmp.out
a write to standard output
before fork
pid = 7156, global = 7, var = 89
before fork
pid = 7155, global = 6, var = 88

OSSI — ver. 1.5 Processes — slide 55

Why two “before fork” messages?

• write() system call not buffered

• write() called before fork(), so one output

• printf() is buffered

∘ line buffered if connected to terminal

∘ fully buffered otherwise; parent and child both have a copy of the unwritten buffer when redirected

• exit() causes both parent and child buffers to flush

OSSI — ver. 1.5 Processes — slide 56

So what does this show?

• It shows that the child is an exact copy of the parent, with all • variable values,

• buffers,

• open files,…

• All are inherited by the child

OSSI — ver. 1.5 Processes — slide 57

Running another program — exec()

• To run another program file

• first call fork() to create a child process

• child calls exec() to replace current copy of parent with a totally

new program in execution

OSSI — ver. 1.5 Processes — slide 58

execve() system call

  #include <unistd.h>

int execve( const char ∗filename, char ∗const argv[], char ∗const envp[] );

• executes the program filename, replaces current process

• Passes the command line in argv[]

• passes the environment variables in envp[]

• Does not return, unless error, when returns with −1

• Usually called through library exec*() calls — see man 3 exec

OSSI — ver. 1.5 Processes — slide 59

fork() — exec() Example

#include <stdio.h>
#include <unistd.h>

int main() {

int pid = fork(); printf( "PID is %d\n", pid ); if ( pid == 0 )

        printf( "I'm the child\n" );

else

        printf( "I'm the parent\n" );

}

OSSI — ver. 1.5 Processes — slide 60

Using execl()
int execl( const char *path,
         const char *arg, ... );

• Parameter number:

1. gives full path of the program file you want to execute

2. gives name of the new process

3. specifies the command line arguments you pass to the program

4. last is a NULL pointer to end the parameter list.

• We must always put a NULL pointer at the end of this list.

OSSI — ver. 1.5 Processes — slide 61

print.c: a program we call

#include <stdio.h>
#include <stdlib.h>

int main( int argc, char ∗argv[] ) {

// argv[0] is the program name int num = atoi( argv[1] ); int loops = atoi( argv[2] ); int i; for ( i = 0; i < loops; ++i )

printf( "%d ", num );

}

OSSI — ver. 1.5 Processes — slide 62

Calling ./print using execl()

#include <stdio.h>
#include <unistd.h>

int main() {

        printf( "hello world\n" );

int pid = fork();

        printf( "fork returned %d\n", pid );

if ( pid == 0 )

               execl( "./print", "print", "1", "100", NULL );

else

               execl( "./print", "print", "2", "100", NULL );

}

vfork() sytem call

• A lightweight fork()

• Designed for running execvp() straight after ∘ modern Linux fork() is very efficient when call exec*()

• Child does not contain an exact copy of parent address space;

• child calls exec() or exit() after fork()

• parent is suspended till child calls fork() or exit()

OSSI — ver. 1.5 Processes — slide 64

wait(), waitpid() system calls
  #include <sys/types.h>
  #include <sys/wait.h>
  pid_t wait( int *status );
  pid_t waitpid( pid_t pid,
                 int *status,
                 int options );

• return process id if ok, 0, or −1 on error

OSSI — ver. 1.5 Processes — slide 65

wait(), waitpid() system calls

• wait() can block caller until child process terminates

• waitpid() has option to prevent blocking

• waitpid() can wait for a specific child instead of the first child • if child has terminated already (it’s a zombie), wait returns im-

mediately, cleaning up the process table data structures for the child

OSSI — ver. 1.5 Processes — slide 66

Part of Simple Shell Program

int main( int argc, char ∗∗argv ) {

char ∗prog name = basename( ∗argv ); print prompt( prog name ); read command(); for ( ;; ) {

int pid = fork(); if ( pid == 0 ) {

execvp( args[ 0 ], args ); } wait( NULL ); print prompt( prog name ); read command();

} }

OSSI — ver. 1.5 Processes — slide 67

Windows and Processes

• Windows provides a Win32 API call to create a process:

CreateProcess()

• Creates a new process, loads program into that process

• CreateProcess() takes ten parameters

OSSI — ver. 1.5 Processes — slide 68

Windows and Processes — 2

• Win32 uses handles for almost all objects such as files, pipes, sockets, processes and events

• handles can be inherited from parent

• No proper parent-child relationship ∘ caller of CreateProcess() could be considered as parent ∘ but child cannot determine it’s parent

OSSI — ver. 1.5 Processes — slide 69

CreateProcess() prototype

• CreateProcess() is much more complicated than

     pid_t fork( void );

• Four of the parameters point to structs, e.g.,

∘ LPSTARTUPINFO points to a struct with 4 members

       ∘ LPPROCESS INFORMATION points to a struct with 18 mem-

bers!

BOOL CreateProcess (

LPCTSTR lpApplicationName, // pointer to executable module

LPTSTR lpCommandLine, // pointer to command line string

LPSECURITY ATTRIBUTES lpProcessAttrib, // process security

LPSECURITY ATTRIBUTES lpThreadAttrib, // thread security

BOOL bInheritHandles, // handle inheritance flag DWORD dwCreationFlags, // creation flags LPVOID lpEnvironment, // pointer to new environment

block

LPCTSTR lpCurrentDirectory, // pointer to current dir name

LPSTARTUPINFO lpStartupInfo, // pointer to STARTUPINFO

LPPROCESS INFORMATION lpProcessInformation // pointer to

//

PROCESS INFORMATION );

CreateProcess()

• Can Specify Program in either 1st or 2nd parameter:

∘ first: location of program to execute

∘ second: command line to execute

• Creation flags:

∘ if 0, runs in existing window

OSSI — ver. 1.5 Processes — slide 71

Example: CreateProcess()

#include <windows.h>
#include <stdio.h>

void main() { STARTUPINFO si; PROCESS INFORMATION pi;

memset( &si, 0, sizeof( si ) );

si.cb = sizeof(si); if ( ! CreateProcess( NULL,

                      "..\\..\\print\\Debug\\print.exe 5 100",

NULL, NULL, TRUE, 0, NULL, NULL, &si,

&pi) )

        fprintf( stderr, "CreateProcess failed with %d\n",

GetLastError() ); WaitForSingleObject( pi.hProcess, INFINITE ); CloseHandle( pi.hProcess ); CloseHandle( pi.hThread );

}

Processes in Linux, Unix, Windows

• Linux often provides 2 or more processes per application

• Example: apache web server parent process watches for connections, one child process per client

• Linux processes have much less overhead than in Windows

• fork() — exec() very

efficient

• posix threads are very efficient, and faster than

fork() — exec()

• Windows have one process per application, but often 2 or more threads

• Windows

CreateProcess()

takes more time than

fork()  —
exec()

•

CreateThread()

takes very much less time than

CreateProcess()

OSSI — ver. 1.5 Processes — slide 73

IPC

Inter Process Communication How Processes can Talk to Each Other

Problem with Processes

• Communication!

• Processes cannot see the same variables

• Must use Inter Process Communication (IPC)

• IPC Techniques include:

∘ pipes, and named pipes (FIFOs) ∘ sockets ∘ messages and message queues ∘ shared memory regions

• All have some overhead

OSSI — ver. 1.5 Processes — slide 75

Interprocess Communication (IPC)

• Pipe — circular buffer, can be written by one process, read by

another

∘ related processes can use unnamed pipes

– used in shell programming, e.g., the vertical bar ‘|’ in

     $ find /etc | xargs file

∘ unrelated processes can use named pipes — sometimes called

fifos

• Messages — posix provides system calls msgsnd() and msgrcv() ∘ message is block of text with a type ∘ each process has a message queue, like a mailbox ∘ processes are suspended when attempt to read from empty

queue, or write to full queue.

OSSI — ver. 1.5 Processes — slide 76

IPC — Shared Memory

• Shared Memory — a Common block of memory shared by many processes

• Fastest way of communicating

• Requires synchronisation (See slide 48)

OSSI — ver. 1.5 Processes — slide 77

IPC — Signals

• Some signals can be generated from the keyboard, i.e., [figure: Control-C] — interrupt (SIGINT); [figure: Control-\]— quit (SIGQUIT), [figure: Control-Z]

  — stop (SIGSTOP)

• A process sends a signal to another process using the kill() system call

• signals are implemented as single bits in a field in the pcb, so cannot be queued

• A process may respond to a signal with:

∘ a default action (usually process terminates)

∘ a signal handler function (see trap in shell programming

notes), or ∘ ignore the signal (unless it is SIGKILL or SIGSTOP)

• A process cannot ignore, or handle a SIGSTOP or a SIGKILL signal.

∘ A KILL signal will always terminate a process (unless it is in

interruptible sleep) ∘ A SIGSTOP signal will always send a process into the stopped

state.

OSSI — ver. 1.5 Processes — slide 78

Signals and the Shell

• We can use the kill built in command to make the kill() system call to send a signal

• A shell script uses the trap built in command to handle a signal

• Ignoring the signals SIGINT, SIGQUIT and SIGTERM:

  trap "" INT QUIT TERM

• Handling the same signals by printing a message then exiting:

  trap "echo 'Got a signal; exiting.';exit 1" INT QUIT TERM

• Handling the same signals with a function call:

  signal_handler() {
      echo "Received a signal; terminating."
      rm -f $temp_file
      exit 1
  }
  trap signal_handler INT QUIT TERM

• Sending a SIGKILL signal to process with pid 3233:

  $ kill -KILL 3233

OSSI — ver. 1.5 Processes — slide 79

Threads Lightweight processes that can talk to each other easily

Threads and Processes

• Threads in a process all share the same address space

• Communication easier

• Overhead less

• Problems of locking and deadlock a major issue • Processes have separate address spaces

• Communication more indirect: ipc (Inter Process Communication)

• Overhead higher

• Less problem with shared resources (since fewer resources to share!)

OSSI — ver. 1.5 Processes — slide 81

Threads have own…

• stack pointer

• register values

• scheduling properties, such as policy or priority

• set of signals they can each block or receive

• own stack data (local variables are local to thread)

OSSI — ver. 1.5 Processes — slide 82

Threads share a lot

• Changes made by one thread to shared system resources (such as closing a file) will be seen by all other threads.

• Two pointers having the same value point to the same data.

• A number of threads can read and write to the same memory

locations, and so you need to explicitly synchronise access

OSSI — ver. 1.5 Processes — slide 83

Threads in Linux, Unix

• posix is a standard for Unix

• Linux implements posix threads

• On Red Hat 8.x, documentation is at

  $ info '(libc) POSIX Threads'

∘ or in Emacs, C-H m libc then middle-click on POSIX

      threads

• Provides:

∘ semaphores, ∘ mutexes and ∘ condition variables

for locking (synchronisation)

OSSI — ver. 1.5 Processes — slide 84

hello.c: a simple threaded program

#include <pthread.h>
#include <stdio.h>

#define NUM THREADS 5 void ∗ print hello( void ∗threadid ) {

printf( "\n%d: Hello World!\n", threadid ); pthread exit( NULL );

} int main() {

pthread t threads[ NUM THREADS ];

int rc, t;

for ( t = 0; t < NUM THREADS; t++ ) {

                printf( "Creating thread %d\n", t );

rc = pthread create( &threads[ t ], NULL, print hello, ( void

∗ ) t );

                if ( rc ) { printf( "ERROR; pthread create() returned

%d\n", rc );

exit( −1 );
}

}

pthread exit( NULL );

}

OSSI — ver. 1.5 Processes — slide 85

How to Compile a POSIX Threads Program

• Need to use the libpthread library

∘ Specify this with the option -lpthread

• Need to tell the other libraries that they should be reentrant (or “thread safe”)

∘ This means that the library uses no static variables that may be overwritten by another thread

∘ Specify this with the option -D REENTRANT

• So, to compile the program ⟨program⟩.c, do:

     $ gcc -D_REENTRANT -lpthread -o ⟨program⟩ ⟨program⟩.c

OSSI — ver. 1.5 Processes — slide 86

pthread create()
  #include <pthread.h>
  void *
  pthread_create( pthread_t  *thread,
                  pthread_attr_t *attr,
                  void *(*start_routine)(void *),
                  void *arg );

• returns: 0 if successfully creates thread

• returns error code otherwise

OSSI — ver. 1.5 Processes — slide 87

pthread create()

• Quite different from fork()

• Thread must always execute a user-defined function

• parameters:

1. pointer to thread identifier

2. attributes for thread, including stack size

3. user function to execute

4. parameter passed to the user function

OSSI — ver. 1.5 Processes — slide 88

Problem with threads:

• Avoid 2 or more threads writing or reading and writing same data at the same time

• Avoid data corruption

• Need to control access to data, devices, files

• Need locking

• Provide three methods of locking:

∘ mutex (mutual exclusion)

∘ semaphores

∘ condition variables

OSSI — ver. 1.5 Processes — slide 89

Race Condition

Race Conditions

• race condition — where outcome of computation depends on sheduling

• an error in coding

• Example: two threads both access same list with code like this:

if ( list.numitems > 0 ) {
     // Oh, dear, better not change to
     // other thread here!
     remove_item( list ); // not here!
     // ...and not here either:
     --list.numitems;
}

OSSI — ver. 1.5 Processes — slide 91

Critical Sections

• critical resource — a device, file or piece of data that cannot be shared

• critical section — part of program only one thread or process should access contains a critical resource

∘ i.e., you lock data, not code

• All the code in the previous slide is a critical section

• Consider the code:

  very_important_count++;

• executed by two threads on a multiprocessor machine (smp =

symmetric multiprocessor)

OSSI — ver. 1.5 Processes — slide 92

Race Condition — one possibility

thread 1  thread 2

read very important count (5)

add 1 (6)

write very important count (6)
                        read very important count (6)

add 1 (7)

                        write very important count (7)

OSSI — ver. 1.5 Processes — slide 93

Example — another possibility

thread 1  thread 2

read very important count (5)
                   read very important count (5)

add 1 (6)

add 1 (6)

write very important count (6)
                   write very important count (6)

OSSI — ver. 1.5 Processes — slide 94

Solution: Synchronisation

• Solution is to recognise critical sections

• use synchronisation, i.e., locking, to make sure only one thread or process can enter critical region at one time.

• Methods of synchronisation include:

∘ file locking ∘ semaphores ∘ monitors ∘ spinlocks ∘ mutexes

OSSI — ver. 1.5 Processes — slide 95

File Locking

• For example, an flock() system call can be used to provide exclusive access to an open file

• The call is atomic

∘ It either:

– completely succeeds in locking access to the file, or

– it fails to lock access to the file, because another thread or process holds the lock

– No “half-locked” state ∘ No race condition

• Alternatives can result in race conditions; for example:

∘ thread/process 1 checks lockfile ∘ thread/process 2 checks lockfile a very short time later ∘ both processes think they have exclusive write access to the

file ∘ file is corrupted by two threads/processes writing to it at the

same time

OSSI — ver. 1.5 Processes — slide 96

Methods of Synchronisation

What is it? mutex, semaphore, condition variables, monitor, spinlock

Synchronisation

• Synchronisation is a facility that enforces ∘ mutual exclusion and ∘ event ordering

• Required when multiple active processes or threads can access shared address spaces or shared i^{/}o resources

• even more critical for smp (Symmetric Multiprocessor) systems ∘ kernel can run on any processor ∘ all processors are of equal importance (there is no one cpu

that is the “boss”) ∘ smp systems include pcs with more than one cpu, as you

might find in the Golden Shopping Centre

OSSI — ver. 1.5 Processes — slide 98

Semaphores

• A variable with three opererations:

∘ initialise to non-negative value ∘ down (or wait) operation:

– decrement variable

– if variable becomes negative, then process or thread executing the down operation is blocked

– has nothing to do with the wait system call for a parent

process to get status of its child ∘ up (or signal) operation:

– increment the semaphore variable;

– if value is not positive, then a process or thread blocked

by a down operation is unblocked.

• A semaphore also has a queue to hold processes or threads waiting

on the semaphore.

OSSI — ver. 1.5 Processes — slide 99

Semaphores — 2

• The up and down semaphore operations are atomic ∘ the up and down operations cannot be interrupted ∘ each routine is a single, indivisible step

• Using semaphores—pseudocode

/∗ only one process can enter critical section at one time: ∗/ semaphore s = 1; down( s ); /∗ critical section ∗/ up( s );

• Initialise semaphore to number of processes allowed into critical

section at one time

OSSI — ver. 1.5 Processes — slide 100

Mutex—posix and Win32 Threads

• mutual exclusion

• Easier to use than semaphores (see slide 50)

• When only one thread or process needs to write to a resource ∘ all other writers refused access

• A special form of the more general semaphore ∘ Can have only two values; ∘ sometimes called binary semaphores.

OSSI — ver. 1.5 Processes — slide 101

mutex — posix Threads Example (1)

• It is good practice to put the mutex together with the data it proects

• I have removed the error checking from this example to save space—in real code, always check library calls for error conditions

#include <pthread.h>
#include <stdio.h>

struct { pthread mutex t mutex; /∗ protects access to value ∗/ int value; /∗ Access protected by mutex ∗/

} data = { PTHREAD MUTEX INITIALIZER, 0 };

OSSI — ver. 1.5 Processes — slide 102

mutex — posix Threads Example (2)

#define NUM THREADS 5

void ∗thread( void ∗t id ) { int i; for ( i = 0; i < 200; ++i ) {

pthread mutex lock( &data.mutex );

++data.value;

               printf( "thread %d:  data value = %d\n",

t id, data.value ); pthread mutex unlock( &data.mutex );

}

pthread exit( NULL );

}

mutex — posix Threads Example (3)

int main() { pthread t threads[ NUM THREADS ]; int rc, t; for ( t = 0; t < NUM THREADS; t++ ) {

            printf( "Creating thread %d\n", t );

pthread create( &threads[ t ], NULL, thread, ( void ∗ ) t );

}

pthread exit( NULL );

}

OSSI — ver. 1.5 Processes — slide 104

posix Condition Variables

• Lets threads sleep till a condition about shared data is true

• Basic operations:

∘ signal the condition (when condition is true)

∘ wait for the condition

– suspend the thread till another thread signals the condition

• Always associated with a mutex

• Very useful

• Missing from Windows:  See http://www.cs.wustl.edu/
  ~schmidt/win32-cv-1.html

OSSI — ver. 1.5 Processes — slide 105

Monitors

• A higher level structure for synchronisation

• Implemented in Java, and some libraries

• main characteristics:

∘ data in monitor is accessible only to procedures in monitor ∘ a process or thread enters monitor by executing one of its

procedures ∘ Only one process or thread may be executing in the monitor

at one time.

• Can implement with mutexes and condition variables.

OSSI — ver. 1.5 Processes — slide 106

Spinlocks

• Used in operating system kernels in smp systems

• Linux uses kernel spinlocks only for smp systems

• a very simple single-holder lock

• if can’t get the spinlock, you keep trying (spinning) until you can.

• Spinlocks are:

∘ very small and fast, and ∘ can be used anywhere

OSSI — ver. 1.5 Processes — slide 107

Summary and References

Summary — Process States, Scheduling

• Scheduler changes processes between ready to run and running

states

∘ context switch: when scheduler changes process or thread • Most processes are blocked, i.e., sleeping: waiting for i^{/}o

∘ understand the process states

∘ why a process moves from one state to another

• Communication between processes is not trivial; ipc methods in-

clude ∘ pipes ∘ messages ∘ shared mem-

ory ∘ signals ∘ semaphores

OSSI — ver. 1.5 Processes — slide 109

Summary — Processes and Threads

• With Linux and Unix, main process system calls are fork(), exec() and wait() — understand the function of each of these

• Windows provides CreateProcess() and various WaitFor...()

Win32 api calls

∘ The WaitFor...() calls have a purpose similar to that of the

wait() system call in Linux and Unix

• Threads are lightweight processes ∘ part of one process ∘ share address space ∘ can share data easily ∘ sharing data requires synchronisation, i.e., locking

OSSI — ver. 1.5 Processes — slide 110

Summary — Synchronisation

• When two threads of execution can both write to same data or

i^{/}o, ∘ Need enforce discipline ∘ Use synchronisation

• We looked at the following methods of synchronisation: ∘ semaphore ∘ mutex ∘ condition variable ∘ monitor ∘ spinlock

• There are other methods we have not examined here.

OSSI — ver. 1.5 Processes — slide 111

References

There are many good sources of information in the library and on the Web about processes and threads. Here are some I recommend:

• A good online tutorial about posix threads: http://www.llnl.gov/computing/tutorials/
  workshops/workshop/pthreads/MAIN.html

• http://www.humanfactor.com/pthreads/ provides links to a lot of information about posix threads

• The best book about posix threads is Programming with POSIX Threads, David Butenhof, Addison-Wesley, May 1997. Even though it was written so long ago, David wrote much of the posix threads standard, so it really is the definitive work. It made me laugh, too!

• Operating Systems: A Modern Perspective: Lab Update, 2nd Edition, Gary Nutt, Addison-Wesley, 2002. A nice text book that emphasises the practical (like I do!)

• Microsoft msdn provides details of Win32 api calls and provides examples of code.

• William Stallings, Operating Systems, Fourth Edition, Prentice Hall, 2001, chapters 3, 4 and 5

• Deitel, Deitel and Choffnes, Operating Systems, Third Edition, Prentice Hall, 2004, ISBN 0-13-1182827-4, chapters 3, 4 and 5

• Paul Rusty Russell, Unreliable Guide To Locking http://kernelnewbies.org/documents/

  kdoc/kernel-locking/lklockingguide.html

• W. Richard Stevens, Advanced Progamming in the UNIX Environment, Addison-Wesley, 1992

• Eric S. Raymond, The Art of UNIX Programming, Addison-Wesley, 2004, ISBN 0-13-

142901-9.

OSSI — ver. 1.5 Processes — slide 112